1. Determine the discharge over an uncontracted horizontal weir that is 1.5 m high and 4.5 m long. The upstream depth is 2.2 m. Determine the Ink height if the same discharge was desired without exceeding an upstream depth of 1.8 m?



2. Determine the discharge over a standard contracted weir when the upstream water depth is 2.1 m. The weir (one end contraction) is 1.8 m high and has a 1.8 m overflow length. Also determine the approach velocity head in the 4-m-wide channel. Is it significant?


1
Expert's answer
2021-12-21T12:24:27-0500

1.

"Q=\frac{Q_0}4=0.0015~\frac{m^3}s,"

"v=\frac QA=\frac Q{\pi r^2}=29.8~\frac ms,"

"\omega=\frac v l=99.5~\frac {rad}s."

2.

"F=pA,"

"A=\pi r^2,"

"p=\frac{\rho v^2}2,"

"F=\frac{\pi r^2\rho v^2}2=855~N."